• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name K. Aniol

    Log entry time 00:02:05 on January13,2005

    Entry number 139108

    keyword=swing shift summary

    16;00, beam is off due to klystron heating problem

    17:40 remote reset of roc4 bib bite high voltage has been fixed but
    remote reset of bigbite fast bus is not yet working as per Jack Segal

    18:00 MCC says they have IOC problems - no estimate of beam now

    18:40 Ran and Konrad went into the Hall to check the PMT associated with p2-y6 on the neutrontdc onlana plots. The pmt produces weak pulse, about 5 to 10 mV, even at 100 volts above the normal operating voltage. The amplifiers and discriminators in the chain from this pmt look ok.

    18:40 MCC starts restoring beam at 4.5uA and no raster.
    18:56 4.5 uA CW in hall

    18:58 run 1753 started, 4.5 uA, LD2, no raster, Kin D1
    19:30 end run 1753, 300K

    19;31 start run 1754, 4.5uA, LD2, no raster, kin D1
    20:13 end run 1754, 300K

    20:14 start run 1755,4.5uA, LD2, no raster, kin D1
    20:50 end run 1755, 300K

    20:52 start run 1756, 4.5uA, LD2,no raster, kin D1
    20:56 end run 1756, 21.6K

    21:00 We ask for 2x2 raster and 10 uA
    HallA esr flow is about 9.3 g/s - with no beam the HPH produce 100W
    The LH2 loop HPH is producing 100W also
    at 10uA the HPH produces 50W and esr flow = 9.3 g/s
    at 10uA the HPH produces 114W and esr flow = 10.4g/s

    21:05 We are informed by MCC that if we want more than 10uA we need to go through special procedures. Contact Bodo and have him talk to the PD.

    21:10 start run 1757, 10uA, 2x2 raster, LD2
    21:20 end run 1757 at 121K

    Contacted MCC to go through procedure for increasing current above 10uA
    21:21 turn HV off neutron detectors -
    21:29 move target to empty position as per MCC.
    21:40 go through on/off procedure with BB current at 10uA
    21:45 No motion of beam spot at beam dump with BB on or off.
    22:00 at 20 uA the HPH on LD2 is producing 67W

    22:10 start run 1758, LD2, 2x2 raster, 25uA
    22:44 end run 1758 with 800K

    22:45 start run 1759, LD2, 2x2 raster, 25 uA
    23:10 veto detectors in upper crate on N-dets have tripped off
    23:17 asked for a controlled access
    23:17 end run 1759 with 800K
    23:31 radcon arrives to escort the controlled access
    23:35 controlled access starts
    23:45 Hall access ends, ask for 5 uA no raster.
    24:00 5 uA on slanted C target